OpenStudy (legends):

Helpppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppp

OpenStudy (legends):

Jack's pet dog has been prescribed 40 mg of a medicine per day. Each pill of the medicine has 12 mg of the medicine. Jack has to give his dog a total of 20 pills. If Jack gave his dog medicine for x days, use the expression below to interpret the situation. : 12 • 20 − 40x Part A: What does 240 represent? Part B: What does the 40x represent? Part C: What does the entire expression represent?

OpenStudy (legends):

@adajiamcneal

OpenStudy (adajiamcneal):

is there like multiple choice? cuz i dont get what theyre asking

OpenStudy (legends):

Nope. These are the questions: Part A: What does 240 represent? Part B: What does the 40x represent? Part C: What does the entire expression represent?

OpenStudy (majesty69):

awww poor alex :(

OpenStudy (adajiamcneal):

pues i dont know what they're asking. ... sorry love

OpenStudy (legends):

Lol it's alright. Thanks anyway.

OpenStudy (here_to_help15):

Each pill has 12 mg of the medicine total of 20 pills 20 mg per pill times 20 pills is 240 mg (of medicine) the total amount of medicine is 240 mg (divided into 20 pills) 40 mg of a medicine per day for x days 40 mg times 1 day = 40 mg 40 mg times 2 days = 80 mg 40 mg times 3 days = 120 mg 12 • 20 − 40x the first part 12*20 = 240 is how much we start with 40x is how much we used up (after x days) 240 mg - 40 mg/day * x days remember that 40x means 40 mg times 1 day = 40 mg 40 mg times 2 days = 80 mg 40 mg times 3 days = 120 mg how much medicine used up. it is so many mg if you start with 240 mg and subtract 40 mg (for example) the answer will be in mg not days. you could look at what you get for x=1, 2, 3 240 - 40*1 = 200 240 - 40*2= 160 240 - 40*3= 120 How much medicine in mg we have left after x days For example, after 3 days, we have 120 mg left (we use the formula 240 - 40*3 = 240-120= 120 mg left) -----__-------------------------------------------- I Part B: 40 x reps that he gives his dog 40 mg of medicine a day Part C: The equation is 12 • 20 − 40x So that means that for each day he gives his dog 20 pills each containing 12 mg for a total of 40 mg a day. Now you do part A :)

OpenStudy (legends):

Thank you soooo much!

OpenStudy (here_to_help15):

Anytime :)